42 students went to youth camp, 56 students went to family camp, 15 went to both camps,how many students went to camp?
Answer by TP(29)

The number of students=56+42-15=98-15=83ANS
Perhaps the best way to see where this answer comes from is by using a VENN DIAGRAM.
Alternatively,
Since 15 went to both camps then 42-15=27 went to the youth camp only.
Also since 15 went to both then 56-15=41 went to the family camp only.
So the total number of students=those who went to youth camp only+those who went to the family camp only+those who went to both=27+41+15=83ANS